# Ground-based Image Deconvolution with Swin Transformer UNet

### v1

@@ 3.3 Test Dataset
We utilized the ESO Distant Cluster Survey, EDisCS (White, S. D. M. et al., 2005), as our benchmark dataset to assess the performance of our deconvolution method. EDisCS is an extensive ESO Large Programme focused on the analysis of galaxy clusters within the redshift range , covering a diverse range of masses with velocity dispersions ranging from approximately km s-1 to km s-1. All clusters benefit from deep , , , and photometry obtained with FORS2 at the VLT. Additionally, a subset of ten clusters was imaged with the ACS at the HST in the F814W filter (Simard et al., 2002). Previous work by Cantale et al. (2016b) employed the deconvolution technique Firedec to analyze spiral disc colours for EDisCS cluster galaxies, investigating trends with cluster masses and lookback-time. In our study, we focus on analyzing a subset of EDisCS clusters at three distinct redshifts—, , and —using our proposed deconvolution method, and go a step further by investigating the star-forming clumps in these galaxies. Table 1 provides a summary of the properties of these clusters and the number of galaxies in each of them.
@@ 4 Results
The EDisCS images in the , , and bands served as inputs for the SUNet deconvolution framework, yielding corresponding deconvolved outputs. Remarkably, the algorithm operates with exceptional speed, requiring only ms to deconvolve a single image on a Titan RTX Turing GPU with 24 GB RAM—approximately times faster than Firedec on average. A histogram of the magnitudes of the same objects in the HST F814W filter is shown in Figure 3. In Section 4.1, we detail our approach to detecting sizes and the number of clumps in the deconvolved outputs. Section 4.2 then presents a comparative analysis of the quality of our SUNet outputs against Firedec (Cantale et al., 2016a). With the validity of our method established, Section 4.3 presents the deconvolved results, followed by a thorough analysis in Section 4.4.
@@ 4.1 Objects Size and Clump Detection
To detect clumps, or small-scale structures, and measure the sizes of galaxies, we employed the SCARLET Python package (Melchior et al., 2018). Utilizing the Starlet transform from SCARLET, our approach involved decomposing images into different scales, where each scale captures a specific frequency component. Ensuring consistency for an unbiased comparison, we maintained fixed algorithm parameters across different bands and objects. All images underwent decomposition into five scales, with the fourth scale chosen for size detection and the second scale for clump detection. A detection threshold was applied to each scale during the process. In the context of clump detection, the Starlet transform was computed using the standard deviation solely within the region enclosed by the size detection outline, rather than considering the entire image. This refined approach ensured more precise thresholding in the Starlet space. Finally, a clump was only considered valid if it lay within the size detection outline, ensuring that background artefacts were excluded. In Figure 4, we present examples of size detection and their corresponding clump detection cases.
@@ 4.2 Comparison with Classical Methods
We conducted a thorough performance comparison between SUNet and Firedec, a classical deconvolution method based on wavelet regularization (Cantale et al., 2016a). For direct comparison with HST quality, we concentrate on the -band outputs for each method, as the EDisCS clusters were exclusively observed in the F814W filter for HST. Both methods exhibit better performance on low-magnitude (or high-SNR) images, with a gradual decline in performance in the high-magnitude (or low-SNR) regime. In this case, the Mean Squared Error (MSE) metric between the deconvolved outputs and the ground truth HST images is not a robust metric for indicating similarity, since it would be biased by the background noise in the HST images (Wang & Bovik, 2009). Instead, we use the Structural Similarity Index (SSIM), a full reference metric which quantifies the similarity between two images by comparing their structural information or spatial inter-dependencies (Wang et al., 2004). An SSIM of 1 implies identical images. The observed trends are illustrated in Figure 5. We further assessed the ability of the deconvolution algorithms to accurately resolve small-scale structures. For this, we leveraged the SCARLET Python package, as detailed in Section 4.1. The fraction of area overlap between small-scale structure detections in the deconvolved outputs and HST is depicted in Figure 6. Based on these metrics, SUNet clearly outperforms Firedec.
Building on the foundation of Firedec, an enhanced method named STARRED was recently introduced by Michalewicz et al. (2023). STARRED brings innovation by incorporating an isotropic wavelet basis known as Starlets (Starck et al., 2015) to refine the regularization process in solving the deconvolution problem. The outputs of Firedec, STARRED, and SUNet are shown in Figure 7. Upon visual inspection of the outputs and the residuals (residual noisy image PSF deconvolved image), it is evident that SUNet consistently generalizes better than Firedec and STARRED.
@@ 4.3 SUNet Deconvolution Results
The methods were put to the test using real ground-based images captured by the FORS2 camera at the Very Large Telescope (VLT) in Chile. Notably, SUNet showed a remarkable ability to generalize effectively to images with entirely different noise properties than those present in the training dataset, as depicted in Figure 8. As illustrated, we are successfully able to recover the morphology and lost small-scale structures. Figure 9 presents the trend in the relative flux error between the SUNet deconvolved outputs and the corresponding HST targets as a function of the total clump size. The uncertainty in the flux level is higher for smaller clumps.
@@ 4.3.1 Resolution Recovery
To gauge the achieved resolution in the deconvolved outputs, we calculated the average ratio between the areas of the smallest detected clump in the SUNet output and its counterpart in the HST image. This ratio, approximately , implies an average SUNet output resolution of around , considering the known HST resolution of .
@@ 4.3.2 False Positives and False Negatives
To assess the reliability of our deconvolution method for real-world applications, we conducted an analysis to estimate the number of false positives and false negatives in our study. We ran SCARLET to detect clumps in both SUNet outputs and HST ground truths. For each HST clump, we checked whether the SUNet-identified clump centroid fell within a -pixel radius of the HST clump centroid. Clumps failing this criterion were considered false positives. The false positive rate was computed across our entire EDisCS dataset by tallying the total count of false positives and dividing it by the overall count of detected clumps in the HST images. However, it is important to note that this result may be biased due to SCARLET’s performance. Instances exist where SCARLET identifies a clump in the SUNet image but misses it in the corresponding HST image, leading to an elevated false positive count. To address this, we employed visual inspection to filter out falsely detected cases. The resultant false positive rate was determined to be approximately . Using a similar approach, we also computed the false negative rate, indicating the probability of missing clumps in the deconvolved images that are present in the HST image. This rate was found to be , signifying a very low probability of missing features.
@@ 4.4 Analysis of Deconvolved EDisCS Galaxies
As a sanity check, we verified that any conclusions drawn regarding the internal properties of our sample galaxies are not influenced by biases in their sizes, either in relation to redshift or disc colours. To this end, the histogram of galaxy sizes grouped by their parent cluster redshift and disc colour is presented in Appendix A. Both plots show that all galaxies have the same global spatial extent.
Following Cantale et al. (2016b), our analysis focuses on the population of galaxies with discs redder than their field counterparts. Some of our sample galaxies have normal colours, hence falling in the colour distribution of the field galaxies, some are bluer (by more than of the colour distribution). A few known physical processes can induce enhanced star-forming activity, but we are instead interested in the possible evidence for quenching mechanisms. Therefore, in the latter case, normal and blue disc galaxies form a common broad class of systems to which we compare the redder ones. Employing the clump detection method outlined in Section 4.1, we computed the histogram of the number of clumps in galaxies, categorized by disc colour, as shown in Figure 10. Note that the only-1-clump case reflects the identification of the central and luminous part (bulge) of the galaxies. As illustrated in Figures 4 and 8, clumps in the -band are brighter than those in the and -bands, in agreement with the spectral energy distribution of young stellar populations. In principle, it is therefore easier to detect clumps in the -band at equivalent photometric depths of the images. This may explain the more continuous distribution in the number of clumps, from to , in the -band. Even so, as witnessed in Figure 10, the general trend is the same from one band to the other. This trend is clear and reveals that the red discs, that were initially identified by Cantale et al. (2016b), have fewer clumps than their bluer counterparts, most likely due to an earlier cessation of star formation. This result opens promising prospects for future studies on larger samples and over larger look-intervals.
@@ 5 Conclusion
We proposed a deconvolution framework involving a two-step process: Tikhonov deconvolution, post-processing with an SUNet denoiser, and an additional debiasing step using multi-resolution support. SUNet was trained on galaxy images from the CANDELS survey and demonstrated superior performance compared to Firedec in the astrophysical context. After establishing the validity of our method, we applied it to deconvolve a set of galaxies from the EDisCS cluster at three different redshifts. Using SCARLET, we provided further analysis of the galaxies in terms of their size and disc colour. We quantified the number of clumps in these galaxies, examining their relationship with disc colour. Our results, based on both quantitative metrics and visual assessments, highlighted the effectiveness of SUNet, showcasing its ability to generalize well to unseen real images with diverse noise properties, which can be attributed to its transformer-based backbone involving the self-attention mechanism.
In summary, this work introduces and evaluates an advanced deconvolution framework applied to ground-based astronomical images. The key findings and contributions include:
Resolution recovery: Based on our SCARLET detection procedure, SUNet demonstrates the capability to recover small-scale structures with an average resolution of approximately and outperforms classical algorithms like STARRED and Firedec (Section 4.3.1).
Generalization to diverse noise properties: The method showcases robust generalization to noise properties different from its training dataset, indicating its adaptability to various observational conditions.
@@
Computational efficiency: The AI-based framework proves to be highly efficient, with an execution time of approximately ms per image, making it around times faster than traditional deconvolution methods like STARRED and Firedec.
Hence, this technique can be used to efficiently identify structures in the distant universe at high spatial resolution from ground-based images. The technique’s applicability to multi-band observations further enhances its utility in studying various astrophysical phenomena. The efficiency of SUNet in processing large datasets and accelerating the deconvolution process opens up opportunities for swift analyses. Access to such a fast and robust deconvolution framework holds the potential to facilitate numerous astrophysical investigations.
6 Reproducible Research
[…]
@@ 6 Reproducible Research
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Acknowledgements.
Appendix A Supplementary Figures
As a validity check, we ensured that any insights into the internal properties of our sample galaxies remain unaffected by size biases, whether concerning redshift or disc colors. This is illustrated in Figure 11, where (a) represents the histogram of galaxy sizes grouped by their parent cluster redshift, and (b) by disc colour. Both plots indicate that all galaxies share the same global spatial extent. Each individual legend in the histograms is normalized such that its probability sums up to one.
@@ Appendix B Impact of debiasing with multi-resolution support on neural networks
For a more rigorous study of the impact of debiasing with multi-resolution support on neural networks, we considered three different neural networks with and without the multi-resolution debiasing: Learnlet (Ramzi et al. 2023), Unet-64 (Ronneberger et al. 2015), and SUNet (Fan et al. 2022). To provide a quantitative comparison, we computed the Structural Similarity Index Measure (SSIM) and the flux error in small-scale structures between galaxies extracted from CANDELS and their simulated degraded versions, as done in Akhaury et al. (2022). The selection of galaxies based on their FWHM and magnitude is depicted in Figure 12. To prevent the background noise in the HST images from biasing our metrics, we fit a Gaussian window around each object with an FWHM equal to its catalogue-derived value. The trends in the metrics, depicted in Figure 13, as a function of the object magnitude, indicate an enhancement in SSIM and flux error for all three networks after debiasing.
@@ Appendix C Hallucinations and the impact of training loss function
To estimate the occurrence of unexpected artefacts or hallucinations introduced by neural networks, we applied the SCARLET detection procedure (as outlined in Section 4.1) with a tight detection threshold to each scale. To determine the number of false positives, we examined whether the centroid of a detection in the neural network’s output fell within a -pixel radius of the HST detection centroid. Detections failing to meet this criterion were considered false positives. Figure 14 illustrates the impact of different loss functions on the hallucination rate as a function of the galaxy FWHM and magnitude for three neural networks: Learnlet, Unet-64, and SUNet. We conducted the experiments on the same test dataset of 2232 galaxies as detailed in Appendix B. Notably, Unet-64 consistently exhibited improved performance when trained with the -loss.

### v2

@@ 3.3 Test dataset
We utilised the ESO Distant Cluster Survey (EDisCS; White, S. D. M. et al., 2005) as our benchmark dataset to assess the performance of our deconvolution method. EDisCS is an extensive ESO Large Programme focused on the analysis of galaxy clusters within the redshift range and covering a diverse range of masses, with velocity dispersions ranging from approximately km s-1 to km s-1. All of the clusters benefited from the deep , , , and photometry obtained with FORS2 at the VLT. Additionally, a subset of ten clusters was imaged with the ACS at the HST in the F814W filter (Simard et al., 2002). Previous work by Cantale et al. (2016b) employed the deconvolution technique Firedec to analyse spiral disc colours for EDisCS cluster galaxies, investigating trends with cluster masses and lookback time. For our study, we focused on analysing a subset of EDisCS clusters at three distinct redshifts—, , and —using our proposed deconvolution method, and we went a step further by investigating the star-forming clumps in these galaxies. Table 1 provides a summary of the properties of these clusters and the number of galaxies in each of them.
@@ 4 Results
The EDisCS images in the -, -, and -bands served as inputs for the SUNet deconvolution framework, yielding corresponding deconvolved outputs. The algorithm operates with exceptional speed, requiring only ms to deconvolve a single image on a Titan RTX Turing GPU with 24 GB RAM—approximately times faster than Firedec, on average. A histogram of the magnitudes of the same objects in the HST F814W filter is shown in Figure 3. In Section 4.1, we detail our approach to detecting sizes and the number of clumps in the deconvolved outputs. Section 4.2 then presents a comparative analysis of the quality of our SUNet outputs against Firedec (Cantale et al., 2016a). With the validity of our method established, Section 4.3 presents the deconvolved results and is followed by a thorough analysis in Section 4.4.
@@ 4.1 Objects size and clump detection
To detect clumps, or small-scale structures, and measure the sizes of galaxies, we employed the SCARLET Python package (Melchior et al., 2018). Utilising the Starlet transform from SCARLET, our approach involved decomposing images into different scales, where each scale captures a specific frequency component. To ensure consistency for an unbiased comparison, we maintained fixed algorithm parameters across different bands and objects. All images underwent decomposition into five scales, with the fourth scale chosen for size detection and the second scale for clump detection. A detection threshold was applied to each scale during the process. In the context of clump detection, the Starlet transform was computed using the standard deviation solely within the region enclosed by the size detection outline, rather than considering the entire image. This refined approach ensured more precise thresholding in the Starlet space. Finally, a clump was only considered valid if it lay within the size detection outline, ensuring that background artefacts were excluded. In Figure 4, we present examples of size detection and their corresponding clump detection cases.
@@ 4.2 Comparison with classical methods
We conducted a thorough performance comparison between SUNet and Firedec, a classical deconvolution method based on wavelet regularisation (Cantale et al., 2016a). For direct comparison with HST quality, we concentrated on the -band outputs for each method, as the EDisCS clusters were exclusively observed in the F814W filter for HST. Both methods exhibited a better performance on low-magnitude (or high-S/N) images, with a gradual decline in performance in the high-magnitude (or low-S/N) regime. In this case, the mean squared error metric between the deconvolved outputs and the ground truth HST images is not a robust metric for indicating similarity since it is biased by the background noise in the HST images (Wang & Bovik, 2009). Instead, we used the structural similarity index measure (SSIM), a full reference metric which quantifies the similarity between two images by comparing their structural information or spatial interdependencies (Wang et al., 2004). An SSIM of one implies identical images. The observed trends are illustrated in Figure 5. We further assessed the ability of the deconvolution algorithms to accurately resolve small-scale structures. For this, we leveraged the SCARLET Python package, as detailed in Section 4.1. The fraction of area overlap between small-scale structure detections in the deconvolved outputs and HST is depicted in Figure 6. Based on these metrics, SUNet clearly outperforms Firedec.
Building on the foundation of Firedec, an enhanced method named STARRED was recently introduced by Michalewicz et al. (2023). STARRED brings innovation by incorporating an isotropic wavelet basis known as Starlets (Starck et al., 2015) that can refine the regularisation process when solving the deconvolution problem. The outputs of Firedec, STARRED, and SUNet are shown in Figure 7. Upon visual inspection of the outputs and the residuals (residual noisy image PSF deconvolved image), it was evident that SUNet consistently generalises better than Firedec and STARRED.
@@ 4.3 SUNet deconvolution results
The methods were put to the test using real ground-based images captured by the FORS2 camera at VLT in Chile. Notably, SUNet showed a remarkable ability to effectively generalise to images with entirely different noise properties than those present in the training dataset, as depicted in Figure 8. As illustrated, we were able to successfully recover the morphology and lost small-scale structures. Figure 9 presents the trend in the relative flux error between the SUNet deconvolved outputs and the corresponding HST targets as a function of the total clump size. The uncertainty in the flux level is higher for smaller clumps.
@@ 4.3.1 Resolution recovery
To gauge the achieved resolution in the deconvolved outputs, we calculated the average ratio between the areas of the smallest detected clump in the SUNet output and its counterpart in the HST image. This ratio, approximately , implies an average SUNet output resolution of around , considering the known HST resolution is .
@@ 4.3.2 False positives and false negatives
To assess the reliability of our deconvolution method for real-world applications, we conducted an analysis to estimate the number of false positives and false negatives in our study. We ran SCARLET to detect clumps in both SUNet outputs and HST ground truths. For each HST clump, we checked whether the SUNet-identified clump centroid fell within a -pixel radius of the HST clump centroid. Clumps failing this criterion were considered false positives. The false positive rate was computed across our entire EDisCS dataset by tallying the total count of false positives and dividing it by the overall count of detected clumps in the HST images. However, it is important to note that this result may be biased due to SCARLET’s performance. Instances exist where SCARLET identifies a clump in the SUNet image but misses it in the corresponding HST image, leading to an elevated false positive count. To address this, we employed visual inspection to filter out falsely detected cases. The resultant false positive rate was determined to be approximately . Using a similar approach, we also computed the false negative rate, indicating the probability of missing clumps in the deconvolved images that are present in the HST image. This rate was found to be , signifying a very low probability of missing features. To obtain a more statistically robust evaluation, we tested the method on another dataset of galaxies extracted from CANDELS (Grogin et al., 2011) and compared it with other neural networks. This work is shown in Appendix C.
@@ 4.4 Analysis of deconvolved EDisCS galaxies
As a sanity check, we verified that any conclusions drawn regarding the internal properties of our sample galaxies are not influenced by biases in their sizes, either in relation to redshift or disc colours. To this end, a histogram of galaxy sizes grouped by their parent cluster redshift and disc colour is presented in Appendix A. Both plots show that all galaxies have the same global spatial extent.
Following Cantale et al. (2016b), our analysis focuses on the population of galaxies with discs redder than their field counterparts. Some of our sample galaxies have normal colours and hence fall into the colour distribution of the field galaxies, and some are bluer (by more than of the colour distribution). A few known physical processes can induce enhanced star-forming activity, but we were instead interested in the possible evidence for quenching mechanisms. Therefore, in the latter case, normal and blue disc galaxies form a common broad class of systems to which we compared the redder ones. Employing the clump detection method outlined in Section 4.1, we computed the histogram of the number of clumps in galaxies, categorised by disc colour, as shown in Figure 10. We note that the only-one-clump case reflects the identification of the central and luminous part (bulge) of the galaxies. As illustrated in Figures 4 and 8, clumps in the -band are brighter than those in the - and -bands, in agreement with the spectral energy distribution of young stellar populations. In principle, it is therefore easier to detect clumps in the -band at equivalent photometric depths of the images. This may explain the more continuous distribution in the number of clumps, from one to six, in the -band. Even so, as we witnessed in Figure 10, the general trend is the same from one band to the other. This trend is clear and reveals that the red discs that were initially identified by Cantale et al. (2016b) have fewer clumps than their bluer counterparts, most likely due to an earlier cessation of star formation. This result opens promising prospects for future studies on larger samples and over larger look intervals.
@@ 5 Conclusion
We have proposed a deconvolution framework involving a two-step process—namely, Tikhonov deconvolution and post-processing with an SUNet denoiser—and an additional debiasing step using multi-resolution support. SUNet was trained on galaxy images from the CANDELS survey and demonstrated superior performance compared to Firedec in the astrophysical context. After establishing the validity of our method, we applied it to deconvolve a set of galaxies from the EDisCS cluster at three different redshifts. Using SCARLET, we provided further analysis of the galaxies in terms of their size and disc colour. We quantified the number of clumps in these galaxies, examining their relationship with disc colour. Our results, based on both quantitative metrics and visual assessments, highlight the effectiveness of SUNet and showcase its ability to generalise unseen real images with diverse noise properties, which can be attributed to its transformer-based backbone involving the self-attention mechanism.
In summary, this work introduces and evaluates an advanced deconvolution framework applied to ground-based astronomical images. The key findings and contributions include the following:
Resolution recovery: Based on our SCARLET detection procedure, SUNet demonstrates the capability to recover small-scale structures, with an average resolution of approximately , and it outperforms classical algorithms such as STARRED and Firedec (Section 4.3.1).
Generalisation to diverse noise properties: The method showcases robust generalisation to noise properties different from its training dataset, indicating its adaptability to various observational conditions.
@@
Computational efficiency: The AI-based framework proves to be highly efficient, with an execution time of approximately ms per image, making it around times faster than traditional deconvolution methods such as STARRED and Firedec.
Our proposed technique can therefore be used with ground-based images to efficiently identify structures in the distant universe at high spatial resolution. The technique’s applicability to multi-band observations further enhances its utility in studying various astrophysical phenomena. The efficiency of SUNet in processing large datasets and accelerating the deconvolution process opens up opportunities for swift analyses. Access to such a fast and robust deconvolution framework holds the potential to facilitate numerous astrophysical investigations.
6 Data availability
For the sake of reproducible research, the codes and the trained models used for this article are publicly available online.
@@ 6 Data availability
The ready-to-use version of our deconvolution method222https://github.com/utsav-akhaury/SUNet/tree/main/Deconvolution.
@@
The repository fork of the SUNet code used for training the network333https://github.com/utsav-akhaury/SUNet.
@@
Link to the trained network weights444https://doi.org/10.5281/zenodo.10287213.
@@
The repository forks of the Learnlet and Unet-64 codes used for comparison in the Appendices B and C555https://github.com/utsav-akhaury/understanding-unets/tree/candels.
@@ Acknowledgements.
Appendix A Supplementary figure
As a validity check, we ensured that any insights into the internal properties of our sample galaxies remain unaffected by size biases, whether concerning redshift or disc colours. This is illustrated in Figure 11, where (a) represents the histogram of galaxy sizes grouped by their parent cluster redshift and (b) represents the same but grouped by disc colour. Both plots indicate that all galaxies share the same global spatial extent. Each individual legend in the histograms has been normalised such that its probability adds up to one.
@@ Appendix B Impact of debiasing with multi-resolution support on neural networks
For a more rigorous study of the impact of debiasing with multi-resolution support on neural networks, we considered three different neural networks with and without the multi-resolution debiasing: Learnlet (Ramzi et al. 2023), Unet-64 (Ronneberger et al. 2015), and SUNet (Fan et al. 2022). To provide a quantitative comparison, we computed the SSIM and the flux error in small-scale structures between galaxies extracted from CANDELS and their simulated degraded versions, as done in Akhaury et al. (2022). The selection of galaxies based on their FWHM and magnitude is depicted in Figure 12. To prevent the background noise in the HST images from biasing our metrics, we fit a Gaussian window around each object with an FWHM equal to its catalogue-derived value. The trends in the metrics, depicted in Figure 13 as a function of the object magnitude, indicate an enhancement in SSIM and flux error for all three networks after debiasing.
@@ Appendix C Hallucinations and the impact of training loss function
To estimate the occurrence of unexpected artefacts or hallucinations introduced by neural networks, we applied the SCARLET detection procedure (as outlined in Section 4.1) with a tight detection threshold to each scale. To determine the number of false positives, we examined whether the centroid of a detection in the neural network’s output fell within a five-pixel radius of the HST detection centroid. Detections failing to meet this criterion were considered false positives. Figure 14 illustrates the impact of different loss functions on the hallucination rate as a function of the galaxy FWHM and magnitude for three neural networks: Learnlet, Unet-64, and SUNet. We conducted the experiments on the same test dataset of 2232 galaxies detailed in Appendix B. Notably, Unet-64 consistently exhibited improved performance when trained with the -loss.